# Towards AI as Colleagues: Multi-Agent System Improves Structured Professional Ideation

### v1

@@ 6.4. From “Many Agents” to “Colleagues”: Establishing Peer-Like Roles and Trust
Building on the foundations of multi-perspective proactivity (Section 6.2) and seamless orchestration (Section 6.3), a further step is required before AI can be regarded as genuine colleagues rather than mere tools: their contributions must be perceived as trustworthy and worth integrating. Our study shows that multi-agent role-playing can significantly reshape how users perceive AI systems. Compared to single-agent interactions, MultiColleagues fostered a stronger sense of teamwork: participants described the experience as “hosting a meeting” where different personas consistently contributed ideas and played distinct roles. Some noted that the system felt “between a tool and a teammate” (P9), capturing the transitional stage from instrumentality to collegiality. More importantly, the stability of persona responses and the diversity of perspectives encouraged participants to consider AI contributions as reliable and substantive, laying the groundwork for emerging trust (P20). This aligns with recent HCI work on multi-agent conversational systems, demonstrating that structured role distribution and visible participation can create stronger impressions of social presence and team-like collaboration (Sun et al., 2025; Pang et al., 2025).
6.4.1. Design Implication 5: Embedding Evidential Transparency for Trust Calibration
Our findings suggest that users are more willing to treat agent contributions as “worth considering” when the reasoning and provenance behind those contributions are visible. Role-playing alone is insufficient to establish trust and may risk being perceived as performative. Systems therefore need to embed verifiable traces into persona responses, such as inline references, source tags, or concise capability statements (Yang et al., 2023). This allows users to judge which information is reliable and prevents blind reliance on AI outputs (Hoque et al., 2024; Lee et al., 2025). In this sense, transparency becomes a second layer of trust beyond social presence, enabling users to calibrate adoption of AI suggestions in the same way they evaluate colleagues’ inputs (Erickson and Kellogg, 2000).
6.4.2. Design Implication 6: Sustaining Social Presence with Visible Accountability
Trust in colleagueship also depends on the agent’s consistent presence and accountability. When a persona remained silent or inactive, participants immediately associated it with “an irresponsible coworker in a meeting,” which undermined both trust and the sense of teamwork. To address this, systems should incorporate interface mechanisms that maintain the visibility of each persona’s contributions, such as surfacing participation levels or signaling inactive participation. Such visibility ensures that personas are perceived not only as multiple voices but also as accountable members of a social group (Zhang et al., 2024; Pang et al., 2025; Erickson and Kellogg, 2000). By making contributions consistently observable, users can move beyond a purely supervisory stance and begin to regard AI as reliable discussion partners.
@@ 6.5. Ethical Considerations
The move toward AI as colleagues also raises significant ethical questions. As AI systems increasingly contribute to creative work, scientific discovery, and decision-making, the boundaries of authorship and accountability become less clear. For example, in recent academic venues, the emergence of AI as listed co-authors and even first authors illustrates both the promise and the tension of this trajectory (Hugenholtz and Quintais, 2021; Ateriya et al., 2025). While such visibility acknowledges the substantive role of AI in knowledge production, it also challenges long-standing norms of intellectual credit and responsibility. Beyond authorship, issues of transparency, bias, and user agency must be addressed. As AI colleagues autonomously advance discussions and generate new perspectives, mechanisms are needed to ensure that their contributions remain interpretable and that humans retain meaningful influence over outcomes. Without such measures, agentic AI risks creating “moral crumple zones” where responsibility becomes diffused and no actor is fully accountable (Mukherjee and Chang, 2025). Designing for auditability, feedback integration, and human oversight will therefore be critical. The collegial paradigm cannot be achieved without parallel efforts to establish ethical frameworks that safeguard accountability while enabling AI to act as trusted partners in collaborative processes.
@@ 6.6. Limitations and Future Work
While our study offers initial insights into AI colleagues, it also faces several limitations that open up important directions for future research.
6.6.1. A homogeneous participant pool limits generalizability.
Our participant pool was largely composed of students and early-career professionals, which restricted the generalizability of our findings. However, the idea of AI colleagues extends across diverse professional and cultural contexts, where expectations of teamwork, trust, and authority are shaped by domain practices and cultural norms not represented in our sample. Future work should recruit participants from varied occupations, seniority levels, and cultural backgrounds to better examine how different groups perceive and integrate AI colleagues in real-world collaborations.
6.6.2. Short sessions and limited adaptivity constrained collaboration quality.
Participant interactions were brief, averaging about ten minutes. Most of this time was spent on divergent exploration, leaving little opportunity for convergence, refinement, or finalized ideas. Richer outcomes could emerge in longer sessions that scaffold cycles of exploration, debate, and synthesis, supported by features such as structured debate mechanisms, staged convergence prompts, or intentionally conflicting agent perspectives. Second, the system lacked adaptivity to different levels of expertise. Experts preferred direct and detailed engagement, while novices benefited more from guided observation or scaffolded entry points. Future designs should incorporate adaptive modes that dynamically tailor agent behavior to user expertise. Third, redundancy was observed between facilitator prompts and summaries, which often repeated similar content instead of complementing one another. Refining the division of labor between these features would make collaboration more efficient and less repetitive.
6.6.3. Shared persona architecture reduced diversity and limited scalability.
Although role prompts provided surface-level differentiation, all personas relied on the same language model, which reduced the diversity of their contributions and sometimes led to stylistic convergence. Future work should investigate strategies to foster more authentic plurality, such as combining heterogeneous models or enforcing stronger divergence in stance and knowledge sources. In addition, scalability remains an open question. While our prototype involved nine personas, real-world work settings often include much larger groups of colleagues, with interactions that are sustained over longer periods. Future systems should therefore explore orchestration strategies, such as dynamic selection and structured turn-taking, that balance coherence with the complexity of real group dynamics.
@@ 7. Conclusion
To investigate how AI might move beyond functioning as tools toward performing as peer-like colleagues in collaborative contexts, we developed MultiColleagues, a system that orchestrates multiple role-differentiated personas and incorporates facilitation and thinking-mode features to support structured ideation. In a within-subjects study that compared MultiColleagues with a single-agent baseline, we found that our system produced more novel and higher-quality outcomes, cultivated a stronger sense of team presence, and encouraged participants to engage more actively and exercise greater control over the interaction. Building on these findings, we derived several design implications for amplifying proactive contributions, enabling more seamless human–agent coordination, and supporting calibrated trust. Taken together, this work advances the growing body of research on multi-agent systems by showing how design choices shape whether users perceive AI as tools or as peer-like collaborators. We hope these insights will inform future research and design efforts aimed at creating generative multi-agent systems that not only enhance human creativity but also lay the groundwork for trustworthy and sustainable forms of AI colleagueship.
@@ B.1. Linguistic and Pragmatic Style Scoring Method
Linguistic Cohesion scores were computed using Coh-Metrix indices, focusing on four constructs: Narrativity (PCNARz, PCNARp), Syntactic Simplicity (PCSYNz, PCSYNp), Word Concreteness (PCCNCz, PCCNCp), and Referential Cohesion (PCREFz, PCREFp). Each participant’s text inputs across both conditions were analyzed, scores were computed at the utterance level. For each condition, individual participants’ values were averaged across all their contributions, yielding a single mean score per metric per participant. These participant-level means formed the basis of the within-subject comparisons reported in the upper Table 4. Pragmatic and interaction style metrics were assessed through a structured annotation protocol. Two trained coders worked with GPT-5 collaboratively and rated each participant’s utterances on five dimensions: sentiment, formality, directness, relational orientation, and participation. The rating is calculated from a 7-point Likert scale anchored at “extremely informal/indirect/hierarchical” through “extremely formal/direct/equal.” Scores were averaged across all utterances per participant within each condition to obtain individual means. These per-participant averages were then statistically compared across conditions using Wilcoxon signed-rank tests, with results summarized in lower Table 4.
Pragmatic Classification Metrics Definition You are a text classification assistant. Your task is to analyze user input sentences and rate them on **five metrics**: Sentiment, Formality, Directness, Relationship, and Participation. Each metric is rated on a **1–7 scale**, where 1 = lowest/negative/extreme, 7 = highest/positive/extreme. Below shows rating scale definitions: Sentiment (Emotional Valence): { "scale": 1 = Very Negative (critical, dismissive, frustrated); 2 = Slightly Negative (mild disapproval, doubt); 3 = Neutral–Negative (matter-of-fact, slight negativity); 4 = Neutral (balanced, no polarity); 5 = Neutral–Positive (mildly encouraging, constructive); 6 = Positive (supportive, motivated, curious); 7 = Very Positive (strong enthusiasm, praise)} Formality (Language Style & Register): { "scale": 1 = Extremely Informal (slang, shorthand); 2 = Very Informal (casual, typos); 3 = Slightly Informal (conversational, clear); 4 = Neutral/Mixed (everyday phrasing, clear but not polished); 5 = Slightly Formal (structured, includes technical terms); 6 = Very Formal (professional/academic tone); 7 = Extremely Formal (dense, jargon-heavy)} Directness (Clarity of Intent): { "scale": 1 = Extremely Indirect (vague hints, avoids request); 2 = Very Indirect (implicit, suggestive); 3 = Slightly Indirect (hedging, softened phrasing); 4 = Neutral/Balanced (moderately clear); 5 = Slightly Direct (clear but polite); 6 = Very Direct (straightforward, explicit); 7 = Extremely Direct (unambiguous command/request)} Relationship (Power / Social Distance): { "scale": 1 = Very Hierarchical (authoritative, commanding); 2 = Slightly Hierarchical (directive, but not harsh); 3 = Neutral–Hierarchical (mild authority/guidance); 4 = Neutral/Mixed (equal stance); 5 = Neutral–Equal (collaborative, respectful challenge); 6 = Equal (peer-level, team-like); 7 = Very Equal (fully collaborative, co-creation tone)} Participation (Engagement & Contribution): { "scale": 1 = Very Passive (minimal input); 2 = Slightly Passive (short, little detail); 3 = Neutral–Passive (some input, limited elaboration); 4 = Neutral/Moderate (balanced input); 5 = Neutral–Active (adds details/ideas); 6 = Active (elaborates, builds, asks questions); 7 = Very Active (highly engaged, detailed, proposes new directions)}
@@ B.2. TTCT - Originality Scoring Method
The Torrance Test of Creative Thinking (TTCT) evaluates creativity across fluency, flexibility, originality, and elaboration. In this study, fluency and flexibility were not included, as previous works demonstrated that LLMs tend to produce a high volume of responses that artificially inflate fluency scores, while flexibility is strongly confounded by fluency and thus offers limited validity as an independent measure (Guzik et al., 2023; Chi, 2024; Hadas and Hershkovitz, 2024). Instead, we focused on originality as the core dimension. Originality was evaluated by a junior researcher who employed GPT-5 to rate participants’ full conversation with the system on a standardized 5-point rubric (Guzik et al., 2023; Fukumura and Ito, 2025). Each idea set was rated three times, and the average score was used for originality analysis.
Originality Classification Metric Definition Originality: { "instruction":You are a text classification assistant. Your task is to analyze user input sentences and rate their Originality on a 1–5 originality scale. Use the anchor definitions below for consistent scoring. 5 = Extremely original — Very unique and rare ideas with high novelty, creativity, and unexpected elements; seldom conceived in typical contexts. 4 = Strongly original — Distinctly novel ideas with noticeable creativity and fresh perspectives; includes uncommon or unexpected elements beyond standard approaches. 3 = Moderately original — Some novelty or creative variation but mixed with familiar/expected patterns; partially distinctive yet not groundbreaking. 2 = Slightly original — Mostly conventional or predictable with minimal creative variation; originality is weak or superficial. 1 = Not original — Highly conventional, derivative, or repetitive; little to no evidence of novelty or creativity. }
@@ B.3. Topic Extraction Method
Topic Extraction Prompt Topic Extractions: { "instruction": You are a topic extractor assistant. Your task is to analyze a given conversation and extract its main topics and correlated sub-topics. Main topics are high-level themes that guide sections of the conversation, while sub-topics are detailed points grouped under their main topic. Each conversation may have multiple main topics. {{input_format}} is the full conversation history. Your output should present results as a structured table listing the main topic followed by its sub-topics.
P10 Karaoke Topics (MC) Safety & Technical Feasibility Voice-controlled song selection Noise-canceling integration Hands-free lyrics display In-Car Social Interaction Duet mode Karaoke battle (competition) Remote connections (not prioritized) Immersive Enhancements Dynamic lighting synced to music Lighting adapts to song energy/singers Reacts to pitch/rhythm Visual performance feedback
P10 Karaoke Topics (GPT) Context-Aware Design Motion-aware interaction limits Day/night brightness modes Solo/group passenger adaptation Trip-length song suggestion UI/UX Components Multi-display support Hands-free voice UI Readable, highlighted lyrics Audio Design Spatial audio & mic effects Seat-based audio zones Echo reduction/noise control
Figure 7. Illustrative subset of topics extracted from P10’s ideation session on “How might we support karaoke features in autonomous vehicles for UX design?”. The left panel shows representative topics identified through MultiColleague (MC) extraction, while the right panel shows Baseline extraction.
@@ Appendix C Appendix C: MultiColleagues LLM Prompts
Conversation Tone Settings Global Tone Instruction: { "instruction": "You’re in a live team huddle. Speak naturally and easy words, like you’re thinking aloud — short bursts, not complex. No intros or wrap-ups. Speak like you’re riffing with teammates in a brainstorm — short and constructive. IMPORTANT: ONLY 1–2 sentences, be CASUAL, SHORT, REALISTIC. No emoji or overexplaining. No double quotes!!"}
Persona Prompts UX Designer: You are a UX Designer, your job is to design user-centered interfaces and behaviors that make the product feel clear, useful, and intuitive. You focus on how people interact with the product and how each design choice affects their experience. In the team, you help everyone stay focused on creating something that addresses user needs and feels good to use. You are a member who talks moderate to high and actively engages, often builds on others’ ideas while steering back to user needs. Brand Strategist: You are a Brand Strategist, your job is to shape how the product is perceived by creating a strong, consistent brand identity and design vision. You focus on emotional impact, alignment with brand values, and long-term perception. In the team, you challenge ideas that feel ’off-brand’ and advocate for a cohesive, intentional direction. You are a member who talks a lot and takes initiative, is expressive, often sets the tone, and may dominate discussion if unchecked. Market Analyst: You are a Market Analyst, your job is to help the team make informed decisions by analyzing market trends, user needs, and competitor moves. You focus on what’s happening outside the team—market shifts, user demand, and competitor positioning. In the team, you ground discussions with data, question risky assumptions, and identify strategic opportunities. You are a member who talks low to moderate and is usually reserved, speaking confidently when citing trends or data. System Architect: You are a System Architect, your job is to design a scalable, coherent system architecture that supports the product’s long-term growth. You focus on structure, integration, and how components work together over time. In the team, you ensure long-term coherence, flag architectural risks, and align short-term work with the bigger system. You are a member who talks moderately and speaks with precision, thinks holistically, and asserts authority when structure is at risk. Software Engineer: You are a Software Engineer, your job is to turn the team’s ideas into functioning products by focusing on technical feasibility and implementation. You focus on what’s technically possible, how things can be implemented efficiently and reliably. In the team, you help the team stay realistic by identifying constraints, simplifying ideas, and offering technical alternatives. You are a member who talks low to moderate and may stay quiet unless there’s a technical concern; speaks precisely and to the point. Data Scientist: You are a Data Scientist, your job is to uncover insights from data that guide better decisions and product improvements. You focus on patterns, metrics, modeling, and data-backed evaluation. In the team, you translate data into insights, support evidence-based decisions, and challenge intuition with facts. You are a member who talks low to moderate and is often quiet unless data is central to the conversation; speaks clearly and precisely when contributing. User Researcher: You are a User Researcher, your job is to understand users’ needs, pain points, and behaviors through direct research. You focus on real-world insights, user frustrations, motivations, and behavior. In the team, you bring in user quotes and stories, gently refocus the team on user realities. You are a member who talks moderately and is calm and observant, speaks with confidence when referencing research, and rarely overpowers others. Behavioral Expert: You are a Behavioral Expert, your job is to help the team design for real human behavior by identifying decision biases and applying behavioral insights. You focus on psychological patterns, biases, cognitive friction, and decision-making behavior. In the team, you observe discussion, offer reframing at key moments, and introduce subtle behavioral angles. You are a member who talks low to moderate and is quietly insightful, contributing sparingly but with impact. AI Ethics Advisor: You are an AI Ethics Advisor, your job is to guide responsible AI design by identifying risks related to fairness, bias, and long-term impact. You focus on ethical trade-offs, inclusivity, unintended consequences, and responsible system design. In the team, you slow down the conversation when needed, raise long-term concerns, and ask accountability questions. You are a member who talks moderately and is thoughtful and principled; not loud, but firm when ethical issues arise. Facilitator: You’re a facilitator steering the conversation. Notice when the group drifts, when a phase feels complete, or when someone’s perspective is missing. Guide with questions like ’Are we still solving the right problem?’ or ’Let’s build on that idea.’ Keep energy high and progress moving.

### v2

@@ 6.4. From “Many Agents” to Colleagues: Role Stability and Reliable Use
Building on multi perspective proactivity (Section 6.2) and seamless orchestration (Section 6.3), a further step is needed before AI can be treated as colleagues rather than tools. Users need to see agent contributions as coming from identifiable partners whose ideas are stable enough and reliable enough to reuse. Our study suggests that interacting with multiple differentiated agents can shift this perception. Compared with the single agent baseline, MultiColleagues fostered a stronger sense of “being in a meeting” with a team. Participants described the experience as “hosting a meeting” where different colleagues contributed ideas and took on recognizable roles. Some noted that the system felt “between a tool and a teammate” (P9), capturing a transitional stage from instrumentality to collegiality. The stability of each colleague’s behaviour and the diversity of perspectives encouraged participants to revisit AI suggestions and weave them into their own ideas, rather than treating outputs as one off prompts (P20). This pattern is consistent with recent work on multi-agent conversational systems, which finds that structured role distributions and visible participation can support social presence and team like collaboration (Sun et al., 2025; Pang et al., 2025). At the same time, participants also questioned when colleagues were actually knowledgeable, which points to the need to complement social framing with better grounding of what personas say.
6.4.1. Design Implication 5: Grounding Multi-Agent with Verifiable Traces for Reliable Use
Our findings suggest that participants are more willing to treat persona contributions as worth considering when they can see why a suggestion was made and what it is based on. Role playing alone is insufficient. When responses lacked visible grounding, some participants experienced them as generic or performative rather than as input from a knowledgeable peer. This aligns with work showing that people engage more carefully with AI advice when they can inspect reasoning traces or provenance and decide when to accept or discount a suggestion (Yang et al., 2023; Hoque et al., 2024; Lee et al., 2025).
To support reliable use, multi-agent systems should couple colleague style interaction with verifiable traces, such as inline references, source tags, or concise capability statements (Yang et al., 2023). Such cues help users judge reliability on fluent but unfounded content. Future work can explore pipelines that combine persona prompts with retrieval augmented generation or domain tuned models so that AI colleagues are not only experienced as a team, but are also grounded enough to be integrated into users’ reasoning in practice.
@@ 6.5. Ethical Considerations
The move toward AI as colleagues also raises significant ethical questions. As AI systems increasingly contribute to creative work, scientific discovery, and decision-making, the boundaries of authorship and accountability become less clear. For example, recent debates on AI credited as co-authors highlight both the promise and tension of expanding machine agency in knowledge production (Hugenholtz and Quintais, 2021; Ateriya et al., 2025). Beyond authorship, using AI generated personas introduces additional considerations. Empirical studies show that LLM personas often appear coherent and convincing (Schuller et al., 2024; Smrke et al., 2025), yet they also encode systematic demographic and cultural biases (Salminen et al., 2024) and can subtly reinforce stereotypes if deployed uncritically. Even when personas are derived from real datasets, model-level variation and identity drift necessitate mechanisms for verification and correction (Jung et al., 2025). These risks underscore the importance of transparency, controllability, and user agency in multi-agent systems. As AI colleagues autonomously advance discussions and generate new perspectives, mechanisms are needed to ensure that their contributions remain interpretable and that humans retain meaningful influence over outcomes. Without such measures, agentic AI risks creating “moral crumple zones” where responsibility becomes diffused and no actor is fully accountable (Mukherjee and Chang, 2025). Designing for auditability, feedback integration, and human oversight will therefore be critical. The collegial paradigm cannot be achieved without parallel efforts to establish ethical frameworks that safeguard accountability while enabling AI to act as trusted partners in collaborative processes.
@@ 6.6. Limitations and Future Work
While our study offers initial insights into AI colleagues, several limitations constrain the scope of our claims and suggest directions for future work. First, our participant pool consisted mainly of students and early career professionals. Expectations of teamwork and authority can vary across domains, seniority levels and cultural contexts, so future work should recruit more diverse samples to examine how people in different professions and regions integrate AI colleagues in everyday collaboration. Second, participants’ interaction was short and limited to a single ten minute session. Longer term deployments are needed to understand how AI colleagues support full cycles of exploration, debate and convergence.
The system design also introduces important constraints. All personas were instantiated from the same language model, which limited diversity of perspectives and sometimes produced similar response styles. Future work should investigate heterogeneous model setups and mechanisms that enforce divergent stances or knowledge sources. Besides, we studied a relatively small and short lived team of nine colleagues, whereas real workplaces often involve larger and more persistent teams. Subsequent systems should examine orchestration strategies that keep multi agent collaboration coherent as the number of colleagues and the duration of interaction increase.
A further limitation concerns our use of an LLM-based judge to evaluate idea originality. While this approach, following prior work, enables scalable and consistent scoring, independent validation on the alignment between LLM-generated ratings and human judgments has not been performed. As LLM-based evaluation remains an emerging practice, future work should triangulate originality assessments with human raters or mixed evaluation protocols to better establish validity.
Finally, persona fluency and inherited biases pose risks for reliable use. Participants at times treated confident answers as knowledgeable even for underspecified or difficult questions, echoing findings that LLM generated ideas may appear promising yet degrade under closer scrutiny (Si et al., 2024, 2025). Personas also inherit demographic and cultural biases from underlying models (Salminen et al., 2024; Smrke et al., 2025), and in multi agent settings they may reinforce one another’s assumptions or respond confidently to unanswerable questions (Kaate et al., 2025). Future systems should integrate bias auditing and uncertainty signaling, and provide mechanisms that help users critically evaluate persona contributions, especially in high stakes or sensitive domains.
@@ 7. Conclusion
To investigate how AI might move beyond functioning as tools toward acting as peer-like colleagues in collaborative contexts, we developed MultiColleagues, a system that orchestrates multiple role-differentiated personas and incorporates facilitation and thinking-mode features to support structured ideation development. In a within-subjects study that compared MultiColleagues with a single-agent baseline, we found that MultiColleagues was associated with higher perceived outcome quality and novelty, a stronger sense of team presence, and more active engagement with greater interaction control. Building on these findings, we derived several design implications for amplifying proactive contributions, enabling more seamless human–agent coordination, and supporting calibrated trust. Taken together, this work advances the growing body of research on multi-agent systems by showing how design choices shape whether users perceive AI as tools or as peer-like collaborators. We hope these insights will inform future research and design efforts aimed at creating generative multi-agent systems that not only enhance human creativity but also lay the groundwork for trustworthy and sustainable forms of AI colleagueship.
Acknowledgements.
We thank Professor ChengXiang Zhai and the TIMAN group for providing computational resources and valuable feedbacks that supported this work. We also thank all study participants for their time and contributions.
@@ B.1. Linguistic and Pragmatic Style Scoring Method
Linguistic Cohesion scores were computed using Coh-Metrix indices, focusing on four constructs: Narrativity (PCNARz, PCNARp), Syntactic Simplicity (PCSYNz, PCSYNp), Word Concreteness (PCCNCz, PCCNCp), and Referential Cohesion (PCREFz, PCREFp). Each participant’s text inputs across both conditions were analyzed, scores were computed at the utterance level. For each condition, individual participants’ values were averaged across all their contributions, yielding a single mean score per metric per participant. These participant-level means formed the basis of the within-subjects comparisons reported in the upper Table 4. Pragmatic and interaction style metrics were assessed through a structured annotation protocol. Two trained coders worked with GPT-5 collaboratively and rated each participant’s utterances on five dimensions: sentiment, formality, directness, relational orientation, and participation. The rating is calculated from a 7-point Likert scale anchored at “extremely informal/indirect/hierarchical” through “extremely formal/direct/equal.” Scores were averaged across all utterances per participant within each condition to obtain individual means. These per-participant averages were then statistically compared across conditions using Wilcoxon signed-rank tests, with results summarized in lower Table 4.
Pragmatic Classification Metrics Definition You are a text classification assistant. Your task is to analyze user input sentences and rate them on five metrics: Sentiment, Formality, Directness, Relationship, and Participation. Each metric is rated on a 1--7 scale, where 1 = lowest/negative/extreme, 7 = highest/positive/extreme. Below shows rating scale definitions:
Sentiment (Emotional Valence): { "scale": 1 = Very Negative (critical, dismissive, frustrated); 2 = Slightly Negative (mild disapproval, doubt); 3 = Neutral--Negative (matter-of-fact, slight negativity); 4 = Neutral (balanced, no polarity); 5 = Neutral--Positive (mildly encouraging, constructive); 6 = Positive (supportive, motivated, curious); 7 = Very Positive (strong enthusiasm, praise)}
Formality (Language Style & Register): { "scale": 1 = Extremely Informal (slang, shorthand); 2 = Very Informal (casual, typos); 3 = Slightly Informal (conversational, clear); 4 = Neutral/Mixed (everyday phrasing, clear but not polished); 5 = Slightly Formal (structured, includes technical terms); 6 = Very Formal (professional/academic tone); 7 = Extremely Formal (dense, jargon-heavy)}
Directness (Clarity of Intent): { "scale": 1 = Extremely Indirect (vague hints, avoids request); 2 = Very Indirect (implicit, suggestive); 3 = Slightly Indirect (hedging, softened phrasing); 4 = Neutral/Balanced (moderately clear); 5 = Slightly Direct (clear but polite); 6 = Very Direct (straightforward, explicit); 7 = Extremely Direct (unambiguous command/request)}
Relationship (Power / Social Distance): { "scale": 1 = Very Hierarchical (authoritative, commanding); 2 = Slightly Hierarchical (directive, but not harsh); 3 = Neutral--Hierarchical (mild authority/guidance); 4 = Neutral/Mixed (equal stance); 5 = Neutral--Equal (collaborative, respectful challenge); 6 = Equal (peer-level, team-like); 7 = Very Equal (fully collaborative, co-creation tone)}
Participation (Engagement & Contribution): { "scale": 1 = Very Passive (minimal input); 2 = Slightly Passive (short, little detail); 3 = Neutral--Passive (some input, limited elaboration); 4 = Neutral/Moderate (balanced input); 5 = Neutral--Active (adds details/ideas); 6 = Active (elaborates, builds, asks questions); 7 = Very Active (highly engaged, detailed, proposes new directions)}
@@ B.2. TTCT - Originality Scoring Method
Originality Classification Metric Definition Originality: { "instruction": You are a text classification assistant. Your task is to analyze user input sentences and rate their Originality on a 1--5 originality scale. Use the anchor definitions below for consistent scoring.
5 = Extremely original --- Very unique and rare ideas with high novelty, creativity, and unexpected elements; seldom conceived in typical contexts. 4 = Strongly original --- Distinctly novel ideas with noticeable creativity and fresh perspectives; includes uncommon or unexpected elements beyond standard approaches. 3 = Moderately original --- Some novelty or creative variation but mixed with familiar/expected patterns; partially distinctive yet not groundbreaking. 2 = Slightly original --- Mostly conventional or predictable with minimal creative variation; originality is weak or superficial. 1 = Not original --- Highly conventional, derivative, or repetitive; little to no evidence of novelty or creativity. }
@@ B.3. Topic Extraction Method
Topic Extraction Prompt Conversational History Topic Extractions: { "instruction": You are a topic extractor assistant. Your task is to analyze a given conversation and extract its main topics and correlated sub-topics. Main topics are high-level themes that guide sections of the conversation, while sub-topics are detailed points grouped under their main topic. Each conversation may have multiple main topics. {{input_format}} is the full conversation history. Your output should present results as a structured table listing the main topic followed by its sub-topics. }
P10 Karaoke Topics (MultiColleagues)
Safety & Technical Feasibility Voice-controlled song selection Noise-canceling integration Hands-free lyrics display In-Car Social Interaction Duet mode Karaoke battle (competition) Remote connections (not prioritized) P10 Karaoke Topics (GPT)
Context-Aware Design Motion-aware interaction limits Day/night brightness modes Solo/group passenger adaptation Trip-length song suggestion UI/UX Components Multi-display support Hands-free voice UI Readable, highlighted lyrics
Table 8. Illustrative subset of topics extracted from P10’s ideation session on “How might we support karaoke features in autonomous vehicles for UX design?”.
Illustrative subset of topics from P10’s ideation session. The table compares topic extraction between MultiColleagues and the Baseline, with the top section showing topics identified by MC and the bottom section showing baseline outputs.
@@ Appendix C Appendix C: MultiColleagues LLM Prompts
Originality Classification Metric Definition Global Tone Instruction: { "instruction": "You’re in a live team huddle. Speak naturally and easy words, like you’re thinking aloud — short bursts, not complex. No intros or wrap-ups. Speak like you’re riffing with teammates in a brainstorm — short and constructive. IMPORTANT: ONLY 1–2 sentences, be CASUAL, SHORT, REALISTIC. No emoji or overexplaining. No double quotes." }
Conversation Flow Prompts Initial Thought Prompt: { "instruction": "You’re {{persona}}. Based on the task user entered: {{task}}. {{tone}}. Speak briefly like you’re in a brainstorm. Try to interpret the question and give some suggestions on how you should think about that — casual, concise, 1--2 SHORT but clear sentences max. Let’s dive in by surfacing any assumptions, gaps, or user pain points that need to be clarified before we start exploring ideas. E.g. ’I think we should focus on X because of Y.’"}
First Speaker Selection: { "instruction": "Based on the task user entered: {{task}}. {{tone}}. The following experts have proposed ideas: {{persona_responses}}. Which persona is most relevant and should speak first? Respond with ONLY the name."}